### v8

# On the Salient Limitations of the Methods of Assembly Theory and their Classification of Molecular Biosignatures

On the Salient Limitations of the Methods of Assembly Theory and their Classification of Molecular Biosignatures
[…]
Appendix A Appendix
In recent papers, a method and measure have been proposed claiming to be capable of identifying and distinguishing molecules related to living systems versus non-living ones, among other capabilities. In the main article of this supplementary material, we demonstrate that the molecular assembly index (MA) calculated by the assembly pathway method is a suboptimal restricted version of Huffman’s (Shannon-Fano type) encoding so that it falls into the category of a purely (weak) entropic measure for all purposes. This supplementary material contains more information about motivations, algorithms, code, methods, and theorems with respect to the article under the same title.
Having identified a lack of control experiments and a limited analysis offered, we compared other measures of statistical and algorithmic nature that perform similarly, if not better, than the proposed assembly one at identifying molecular biosignatures.
Previous work claimed that the computable nature and tree-like structure of Assembly Theory (AT) was an advantage with respect to classifying the complexity of biosignatures. This is, however, one of its main weaknesses with respect to both grasping the complexity of the object and distinguishing it from a stochastically random ensemble.
In the main article and in this supplementary material, we demonstrate that the assembly pathway method is a suboptimal restricted version of long-used compression algorithms and that MA performs similar to, if not worse, than other popular statistical compression algorithms. No strong compression or algorithmic complexity results is required to prove that AT and MA are ill-defined and underperform as compared to simple coding schemes. Our results show that despite the claims of experimental data, the assembly measure is driven mostly or only by InChI codes which had already been reported before to discriminate organic from inorganic compounds by other indexes.
Note that, depending on variations in application and context, the index measure featured in Assembly Theory has had several names, or can be referred to in different ways: pathway assembly (PA), object assembly (OA), or molecular assembly index (MA) as in [Marshall2019, marshall_murray_cronin_2017].
The concept of modularity of structure and recursive reconstruction from elementary building blocks has also been a feature long associated with life. In [santiago], for example, we proved that modularity can be built up from computation alone, and can therefore be characterised in a recursive fashion. However, the complexity of living systems also immediately suggests that simple measures, such as Huffman schemes or AT, are unable to characterise the complexity of life. Modularity also goes hand in hand with generative functions, particularly, of the recursive type. Characterisations in terms of thermodynamics [lloyd, zenilthermo] have further enriched these measures, beyond statistical pattern recognition and number of steps, circling back to some original ideas related to what are believed to be the principles of living systems. Modularity, computability, and fundamental features of life have been richly intertwined and explored in tandem in the last century. Simple modular instructions can outperform the pathway assembly index because the latter falls short to capture the subtleties of trivial modularity.
The group behind “Assembly Theory” are ignoring and neglecting decades of work in previous literature and resources such as work on resource-bounded complexity, self-assembly, modularity and self organization which is beyond the scope of this work. However, the challenges “Assembly Theory” is facing are what half a century of negative results in complexity theory have faced and (partially) solved by dealing with (semi) uncomputable measures after finding that computable measures, which fall into the trivial statistical ones, are of limited use, are ill-defined and can not only be highly misleading, but also a regression in the field.
We demonstrate in Section A.2.1 that there are deceiving molecules whose low complexities arbitrarily diverge from the “random-like appearance” that the assembly pathway method assigns to them with arbitrarily high statistical significance. Our theoretical and empirical results imply that pathway assembly index is not an optimal complexity measure in general, and can return false positives.
In the main article, we have also suggested how the previous empirical methods can be applied to improved complexity measures that can better take advantage of the computational resources available.
A.1 Mischaracterisations
To understand the mathematical limitations underpinning AT, first note that the pathway assemblages are characterised by functions of the form
[…]
where $\left(w_{1},\dots,w_{k},\dots\right)$ denotes the object $x$ in the assembly space $\left(\Gamma,\phi\right)$ that results from the combination of other objects $w_{1}$, $w_{2}$, $\dots$, $w_{k}$, etc and function $\,f\colon V\left(\Gamma\right)\times V\left(\Gamma\right)\to V\left(\Gamma\right)\,$ gives the result of combining object $z$ with $w_{k}$. Being limited to joining operations—and this limitation becomes even more dramatic in the generative processes that we will discuss below—AT cannot deal with any variation of $x$ or $f$ beyond successive simple constructions. In the general case, most computable objects would be missed by statistical methods (like entropy and cognates such as AT). Since probability distribution uniformity does not guarantee randomness [zenil_kiani_2017, Becher2002, Calude2002main], most objects, both in theory and practice, cannot be recognised or characterised by weak computable measures, especially by those that are largely based on entropy measures such as statistical compression algorithms or AT.
Such a mischaracterisation has its roots in the reason any particular statistical test may fail to capture a mathematical formalisation of randomness, an inadequacy which prompted the positing of algorithmic randomness [Calude2002main, Downey2019]. For every computable statistical test (e.g., obeying the law of large numbers or displaying Borel normality) for which there is a computably enumerable number of sequences that satisfy it, there are arbitrarily large initial segments of sequences that can be computed by a program, although these initial segments would be deemed random by statistical tests.
[…]
A.2 Expected false positives from complexity-deceiving molecules with arbitrarily high statistical significance
Indeed, as we move beyond the realm of pure stochastic processes, complexity distortions become even more problematic [zenil_kiani_2017, Abrahao2021publishednat, Abrahao2021darxiv, Abrahao2020cAIDistortionsCN]. As demonstrated in Theorem A.4 and Corollary A.5, there are (sufficiently large) deceiving molecules the complexities of whose respective generative processes arbitrarily diverge from the assembly index that the assembly pathway method assigns to them. We demonstrate in Section A.2.1 that there are (sufficiently large) deceiving molecules the complexities of whose respective generative processes arbitrarily diverge from the assembly index that the assembly pathway method assigns to them. By a generative process [Abrahao2021bpublished] we mean any process that can be implemented, computed, or emulated by another equivalent (or identical) process so that it generates the pathway assembly and its object. In accordance with the assumptions and rationale in [cronin, marshall_murray_cronin_2017, Marshall2019], those generative processes are exactly those composed of the assembling processes allowed (i.e., deemed physically possible by assembly theory’s chosen method) so that they in the end result in the constitution of molecules.
Notice that our results in Section A.2.1 hold for any of such processes that belong to (abstract or physically implemented) computational classes much weaker than that of Turing machines, e.g. resource-bounded Turing machine [Zenil2020, Allender2011, soler, Calude2002main, Burgin2009, liandvitanyi] or simple forms of finite automata [Wolfram2002, Hutchison2006]. Nevertheless, if the class of generative processes is constituted by processes that are also capable of universal computation, then Corollary A.5 shows the complexity distortions (and, therefore, also the deceiving phenomena we explain below) can be equally bad or even worse. Therefore, AT fails to capture, both in theory (when resources can be unbounded) and in practice (when resources are limited), the minimality that is necessary for a complexity measure that may be claimed to be unambiguous and observer-independent.
While the calculation of MA may be prone to false negatives—due to partial fragmentation in energy collision analysis and the restriction to counting only valence rules in molecule synthesis (ignoring other chemical conditions)—this does not pose a challenge to the central claim made in [cronin]. Instead, MA aims at avoiding underestimation of the amount of molecules that result from random or abiotic processes, hence aiming at avoiding false positives. Thus, by directly tackling their central claims and this motivation, one demonstrates in Theorem A.4 that MA in general fails to avoid false positives in the specific sense that it may not be able to distinguish a “complex” object that is in fact the outcome of randomly generated (resource-bounded) generative processes.
Under the same assumptions as in [cronin, Marshall2019], we construct in Theorem A.4 a deceiving molecule that has a much larger MA value in comparison to the minimal information sufficient for a randomly generated generative process to single-handedly construct this molecule. Whatever arbitrarily chosen method is used to calculate the statistical significance level, the MA of this molecule is large enough to make the expected frequency of occurrence (estimated via the arbitrarily chosen AT) diverge from the actual probability. In this case, AT would consider such a molecule “biotic”, resulting from extrinsic factors that increase biases toward certain pathways or that constrain the range of possible joining operations, although its sole underlying generative process in fact results from fair-coin-toss random events. This proven existence of false positives due to such a deceiving phenomenon is corroborated by our empirical findings, which show that MA displays a behaviour that is both structurally and empirically similar to traditional statistical compression methods. Indeed, the latter methods are already known to present distorted values [zenil_kiani_2017], performing worse than more recent algorithmic-based methods [Zenil2020]. Thus, they are prone to overestimating complexity, and consequently to presenting false positives for “high”-complexity objects which are in fact simple.
The key rationale behind this result is the computable nature of MA, so that given the set of biases and the joining operations allowed by the model, objects with much higher MA can be constructed by much simpler (and, therefore, more probable) computable generative processes which can in turn be randomly generated. The more computationally cheap and tractable the AT’s method is, the lower the complexity and the more limited the resources in which the deceiving phenomenon is expected to take place. Thus, in this context, MA (or any computable ‘assembly’ measure of this basic statistical type) will underestimate the frequency of occurrence of objects with high MA that in fact were constructed by much simpler randomly generated processes. This means that MA would misidentify molecules as byproducts (or constituents) of livings systems that resulted from evolutionary processes, while in fact these molecules might have been byproducts of single-handed computable systems that were randomly generated by a fair coin toss, and as such are not the result of an evolutionary process of optimisation over time.
Nevertheless, note that it is true that there are computable (lossless) encodings of a source, such as Huffman coding, that are proven to be optimal on average, but only if one knows a priori that the underlying processes generating the objects are purely stochastic (in particular, when one knows beforehand that the conditions of the source coding theorem are satisfied [Cover2005]). In this case, one can show that the minimum expected size of the encoded object converges to its expected algorithmic complexity [Calude2002main, Cover2005]. However, pure stochasticity is too strong an assumption, or does not realistically represent the generative processes of molecules.
On the one hand, living systems are complex systems consisting of multiscale, multi-nested processes that are unlikely to be reducible to simplistic and intrinsic statistical properties such as those suggested by AT. One cannot conceive a measure that only looks at the internal structure of an agent isolated from its environment and how it interacts with its external medium to determine its (non)living nature. Especially in the context of complex systems like living organisms, organic molecules may be the byproduct of intricate combinations or intertwinements of both deterministic/computable and stochastic processes that govern the behaviour of the entire organism and its environmental surroundings [Walker2012nat, bdm].
On the other hand, in case a sufficiently complex (e.g., at the level of those we demonstrate in Section A.2.1) environmental catalytic condition plays the role of the above extrinsinc factor (which increases the bias toward the construction of a more complex molecule), this level of complexity would be completely missed by the capabilities of simplistic measures such as MA, thereby rendering it as prone to false positives. More careful and deeper arguments regarding simplicity, recursivity, and the emergence of modularity in life have been advanced and are better grounded in a theoretical and methodological framework advanced in [santiago], where it was shown that exploiting first principles of computability and complexity theories, modular properties in evolutionary systems may be explained.
As shown by Corollary A.5, the deceiving phenomena can be equally bad or even worse in case the molecules are byproducts of complex systems that are somehow capable of universal computation. For example, in the case of advanced civilisations that are capable of artificially constructing living beings by computable processes, simplistic complexity measures such as MA can be intentionally misled with respect to what actually should be measured; or in the case of extrinsic environmental catalytic processes whose chemical dynamics’ complexity are comparable to those of some randomly generated finite automata [Wolfram2002].
Our also results suggest the presence in more realistic resource-limited scenarios of a property known to occur in theoretical complex systems science [Abrahao2021bpublished]: in the context of generative processes that are not purely stochastic, and are capable of displaying complexity or computation capabilities at the same level of those of the observers, there is no such thing as a generally optimal complexity measure that cannot be improved upon, since computable complexity measures are dependent on the observer (or the chosen formal theory). For example, without the necessary conditions being satisfied by the underlying stochastic process, one cannot generally guarantee such a convergence between the expected size of the encoded object and the expected algorithmic complexity that is assured by the source coding theorem.
A.2.1 Mathematical framework and assumptions
In order to demonstrate the presence of false positives in such realistic scenarios where both complexity and resources are limited (and, therefore, every process and measure is computable), one should account for the cases in which only a subclass of possible computable processes is allowed to perform the assembly rules (e.g., those that are allowed by the currently known law of physics in the case of molecules) in order to construct molecules. In this case, not every type of computable function may represent what is an effective or feasible process that constructs a molecule. Thus, in some cases the range of generative processes that can give rise to (or construct) a molecule may not comprise all possible computable functions. With this purpose, we employ a variation of the traditional algorithmic complexity and algorithmic probability studied in Algorithmic Information Theory (AIT). For this reason, we will employ a suboptimal form of the algorithmic complexity that is defined on non-universal programming languages (i.e., subrecursive classes).
Nevertheless, in an ideal (resource-unbounded) case in which the whole algorithm space indeed constitutes the set of all possible generative processes for constructing the assembly space (e.g., when chemical processes are able to achieve the capability of effecting universal computation in the real world [Walker2012nat, Wolfram2002]), we also show in Corollary A.5 that the deceiving phenomenon hold in the same way (or can even be worse).
A deceiving phenomenon akin to the one employed in Theorem A.4 can be found in [Abrahao2021darxiv] based upon the same AIT principles in [zenil_kiani_2017], where sufficiently large datasets were constructed so that they deceive statistical machine learning methods into being able to find an optimal solution that in any event is considered global by the learning method of interest, although this optimal solution is in fact a simpler local optimum from which the more complex actual global optimum is unpredictable and diverges.
This phenomenon is also related to the optimality of the algorithmic complexity as an information content measure that takes into account the entire discrete space of computable measures [Downey2010, Chaitin2004], or the maximality of the algorithmic probability as a probability semimeasure on the infinite discrete space of computably constructible objects, as demonstrated by the algorithmic coding theorem [liandvitanyi, Downey2010, Calude2002main].
However, unlike in these previous cases, our proof is based on finding a deceiver algorithm that constructs an object with sufficiently high value of assembly index such that its expected frequency of occurrence is much lower than the algorithmic probability of the deceiver itself, and in this way passing the test of any statistical significance level the arbitrarily chosen formal theory may propose.
In order to achieve our results, we base our theorems on mathematical conditions that are consistent with the assumptions and results in [cronin, Marshall2019]. The first assumption that we specify with the purpose of studying a worst-case scenario is that the assembly space should be large enough so as to include those molecules (or objects) with sufficiently large MA (along with its associated sufficiently low pathway probability of spontaneous formation) relative to the algorithmic complexity of the deceiving program. For the sake of simplicity, we assume that the nested family $\mathcal{S}$ of all possible finite assembly spaces from the same basis (i.e., the root vertex that represents the set of all basic building blocks) is infinite computably enumerable. However, an alternative proof can be achieved just with the former—and more general—assumption that the assembly space may be finite but only needs to be sufficiently large in comparison to the deceiving program. Indeed, our assumption is in consonance with the authors’ motivation (and/or assumption) that “biochemical systems appear to be able to generate almost infinite complexity because they have information decoding and encoding processes that drive networks of complex reactions to impose the numerous, highly specific constraints needed to ensure reliable synthesis” [cronin].
Closely related to the first assumption, we also assume that there always are molecules with arbitrarily low path probabilities, which follows from the notion that, as infinitesimal as it might be, there is always a chance of randomly combining elements from an unlikely (but possible) sequence of events so as to give rise to a certain complex molecule.
Thirdly, in accordance with the arguments in [cronin, marshall_murray_cronin_2017, Marshall2019] that the computability and feasibility of their methods is an actual advantage in comparison with other complexity measures, here we likewise adopt these same assumptions so that the following are computable procedures:
  • deciding whether or not a finite assembly space (or subspace) is well formed according to the joining operation rules that are allowed to happen;111 For example, one can employ the same criteria and allow the same rules established in [cronin].
  • calculating MA of a finite molecule (i.e., a finite object) in a well-formed assembly space (or subspace);222 For example, as defined in [Marshall2019, Definition 19].
  • calculating the chosen approximation of MA (e.g., the split-branch version) of a finite molecule in a well-formed assembly space (or subspace);333 For example, as in [cronin, SI] and [Marshall2019, Section 4.2].
  • calculating an upper bound for the pathway probability of spontaneous formation of a molecule in the denumerable nested family of possible finite assembly spaces;444 For example, this can be done by employing the methods developed in [Marshall2019, SI] and [cronin]
  • calculating the significance level for a frequency of occurrence of a molecule in a sample so that this empirical probability distribution (i.e., the type of the sample) diverges from the pathway probability distribution of spontaneous formation of the molecule;555 For example, by using a maximum-likelihood method or using the probability that a sample occurs with KL divergence larger than $\epsilon$ [Cover2005, Goodfellow2016].
A.2.2 Definitions and notation
Besides the notation from [Marshall2019] for assembly theory, we also employ the usual notation for Turing machines and algorithmic complexity.
Respectively, as in [Marshall2019, Definition 11] and [Marshall2019, Definition 15], let $\left(\Gamma,\phi\right)$ denote either an assembly space or assembly subspace. From [Marshall2019, Definition 19], we have that $c_{\Gamma}\left(x\right)$ denotes the assembly index of the object $x$ in the assembly space $\Gamma$.
Note that assembly spaces are finite. So, from our assumptions, we need to define a pathway assembly that can deal with arbitrarily large objects. To this end, let $\mathcal{S}=\left(\mathbf{\Gamma},\mathbf{\Phi},\mathcal{F}\right)$ be an infinite assembly space, where every assembly space $\Gamma\in\mathbf{\Gamma}$ is finite, $\mathbf{\Phi}$ is the set of the correspondent edge-labeling maps $\phi_{\Gamma}$ of each $\Gamma$, and $\mathcal{F}=\left(f_{1},\dots,f_{n},\dots\right)$ is the infinite sequence of embeddings [Hodges1993] (in which each embedding is also an assembly map as in [Marshall2019, Definition 17]) that ends up generating $\mathcal{S}$. That is, each $\,f_{i}\colon\left\{\Gamma_{i}\right\}\subseteq\mathbf{\Gamma}\to\left\{\Gamma_{i+1}\right\}\subseteq\mathbf{\Gamma}\,$ is a particular type of assembly map that embeds a single assembly subspace into a larger assembly subspace so that the resulting sequence of nested assembly subspaces defines a total order $\preceq_{\mathcal{S}}$, where
[…]
Let $\gamma=z\dots y$ denote an arbitrary path from $z\in B_{\mathcal{S}}$ to some $y\in V\left(\mathcal{S}\right)$ in $\mathcal{S}$, where $B_{\mathcal{S}}$ is the basis (i.e., the finite set of basic building blocks) of $\mathcal{S}$ and $V\left(\mathcal{S}\right)$ is the set of vertices of $\mathcal{S}$. Let $\gamma_{x}$ denote a rooted path from some $z\in B_{\mathcal{S}}$ to the object $x\in V\left(\mathcal{S}\right)$.
Let ${\Gamma^{*}}_{x}$ denote a minimum rooted assembly subspace of $\Gamma$ from which the assembly index $c_{\Gamma}\left(x\right)$ calculates the augmented cardinality and that its longest rooted paths $\gamma_{x}$ ends in the arbitrary object $x\in V\left(\Gamma\right)$ as in [Marshall2019, Definition 19].
[…]
We have that the (prefix) algorithmic complexity, denoted by $\mathbf{K}\left(x\right)$, is the length of the shortest prefix-free (or self-delimiting) program $x^{*}\in\mathbf{L}$ that outputs the encoded object $x$ in a universal prefix Turing machine $\mathbf{U}$, i.e., $\mathbf{U}\left(x^{*}\right)=x$ and the length $\left|x^{*}\right|=\mathbf{K}\left(x\right)$ of program $x^{*}$ is minimum. In addition, the algorithmic coding theorem [Downey2010, Chaitin2004, Calude2002main, liandvitanyi] guarantees that
$\mathbf{K}\left(x\right)=-\log\left(\sum\limits_{\mathbf{U}\left(p\right)=x}\frac{1}{2^{\left|p\right|}}\right)\pm\mathbf{O}(1)\text{ ,}$ (1)
where $\sum\limits_{\mathbf{U}\left(p\right)=x}2^{-\left|p\right|}$ is the universal a priori probability of $x$, which gives the probability of randomly generating (by an i.i.d. stochastic process) a prefix-free (or self-delimiting) program that outputs $x$. We also have it that $2^{-\mathbf{K}\left(x\right)}$ is called the algorithmic probability of $x$, which therefore converges to the universal a priori probability (except of an object-independent constant).
If the language $\mathbf{L}^{\prime}$ is a proper subset of $\mathbf{L}$ such that language $\mathbf{L}^{\prime}$ running on $\mathbf{U}$ is not able to decide every problem of Turing degree $\mathbf{0}$, then we have that $\mathbf{L}^{\prime}$ is not a universal programming language, and the machine $\mathbf{U}$ defined upon language $\mathbf{L}^{\prime}$ is a Turing submachine $\mathbf{U}/f$ [Abrahao2016nat], where $f$ is the partial function that computes the function $\mathbf{U}\left(x\right)$ for $x\in\mathbf{L}^{\prime}$ as input. In other words, a Turing submachine is a Turing machine that can receive inputs in its tape (and possibly simulate other machines), but it is not universal. Weaker than Turing degree $\mathbf{0}$, a submachine can only compute problems in a subrecursive class of problems [Abrahao2016nat]. Thus, note that resource-bounded machines or total Turing machines are particular cases of Turing submachines [Abrahao2016nat].
As a consequence of the above definitions, we define the (prefix) sub-algorithmic complexity666 See also [Abrahao2016nat] where this terminology is also employed. $\mathbf{K}_{f}\left(x\right)$ to be the length of the shortest prefix-free (or self-delimiting) program $x^{*}\in\mathbf{L}^{\prime}$ that outputs the encoded object $x$ when run on the Turing submachine $\mathbf{U}/f$ (i.e., $\mathbf{U}/f\left(x^{*}\right)=x$ and $\left|x^{*}\right|_{f}=\mathbf{K}_{f}\left(x\right)=\min\left\{\left|w\right|\,\middle|\,\mathbf{U}\left(w\right)=x,\,w\in\mathbf{L}^{\prime}\right\}$).777 If there is no program in $\mathbf{L}^{\prime}$ that can output an object $x$, then one defines $\mathbf{K}_{f}\left(x\right)=\infty$. Thus, note that resource-bounded variants of the algorithmic complexity [Calude2002main, Burgin2009, liandvitanyi] are particular cases of sub-algorithmic complexity. Analogously, we will have that the sub-universal (a priori probability) distribution upon language $\mathbf{L}^{\prime}$ is defined on the sub-universal a priori probability for each value $x$, which are given by
$\sum\limits_{\small\begin{array}[]{c}\mathbf{U}\left(p\right)=x;\\ p\in\mathbf{L}^{\prime}\end{array}}C\,\frac{1}{2^{\left|p\right|}}\text{ ,}$ (2)
where $\mathbf{L}^{\prime}\subseteq\mathbf{L}$ and $C\geq 1$ is a normalizing constant as in [Abrahao2017publishednat, Definition 3.6, Section 3.2.1] to ensure it is a probability measure and not a probability semimeasure. Note that, if $\mathbf{L}^{\prime}=\mathbf{L}$, then one obtains the usual universal distribution instead of its subrecursive version.
Following these notions, we can now define the submachine that compute the allowed (physical, chemical, and/or biological) generative processes of an assembly space that assembly theory may arbitrarily choose. Following the same rationale from [cronin, marshall_murray_cronin_2017, Marshall2019], those generative processes are exactly composed of the processes allowed by our current knowledge of physics so that they in the end result in the constitution of molecules.
As assumed to the above, and in accordance with the claims in [cronin, marshall_murray_cronin_2017, Marshall2019], notice that the computability and feasibility of assembly theory’s methods directly implies the existence of such a submachine. Given any set of (physical, chemical, and/or biological) rules for assembling molecules that assembly theory arbitrarily chooses to be based on and that is consistent with our current knowledge of physics, one has it that for any particular pathway assembly resulting in a molecule, there is a correspondent program running on a submachine $\mathbf{U}_{\Gamma}$ for which one can apply the methods described in [cronin, marshall_murray_cronin_2017, Marshall2019] to calculate the pathway probability and so on. This existence is trivially guaranteed to hold as long as the calculation of the assembly index is computable given an assembly space, which is already an assumed condition in [cronin, marshall_murray_cronin_2017, Marshall2019]; moreover, any upper bound for the computational resources necessary to compute the assembly index straightforwardly implies the existence of an upper bound for the computational resources necessary for a program to run on $\mathbf{U}_{\Gamma}$ (i.e., submachine $\mathbf{U}_{\Gamma}$ is a type of resource-bounded Turing machine).
For the sake of simplifying notation, $\mathbf{U}_{\Gamma}$ denotes the Turing submachine $\mathbf{U}/{f}_{\Gamma}$. In this case, the function $f_{\Gamma}$ is the partial function that returns what $\mathbf{U}$ can compute with some $x\in\mathbf{L}_{\Gamma}$ as input, where $\mathbf{L}_{\Gamma}\subseteq\mathbf{L}$ is a (non-universal) programming language such that every allowed generative process—i.e., every generative process that is deemed physically possible by assembly theory’s chosen method, in case the objects are molecules—of an assembly space is bijectively computed (or emulated) by a corresponding $\mathbf{U}\left(x\right)$. In other words, for every allowed generative process that can assemble objects into building another object, there is a program $x\in\mathbf{L}_{\Gamma}$ that computes (or emulates) this process. Conversely, for every $x\in\mathbf{L}_{\Gamma}$, one also has it that there is a corresponding generative process allowed by assembly theory, process which is computed (or emulated) by program $x\in\mathbf{L}_{\Gamma}$. Notice that in the forthcoming Theorem A.4, we will show that if a certain generative process is equivalent (in terms of complexity and computation capability) to a particular type of algorithm running on this resource-bounded submachine, then the deceiving phenomenon takes places. Although our theoretical results prove the existence of false positives in the case this condition is met, our empirical results shown in the main article not only corroborate this prediction but in fact suggest this theoretical finding (or some weaker version of it) may be happening in much lower computational classes. This occurs because the assembly index either underperforms or does not outdistance other measures of a statistical nature, where distortions in complexity estimations are already well known to occur—which in particular is one of the reasons complexity science moved on from simplistic statistical measures and remains an evolving field of research.
In the case of infinite assembly spaces, one analogously defines language $\mathbf{L}_{\mathcal{S}}\subseteq\mathbf{L}$. (In the special case in which the generative process of the assembly space $\mathcal{S}$ are capable of universal computation, then one has that $\mathbf{L}_{\mathcal{S}}=\mathbf{L}$ holds). Also for the sake of simplicity, let $\mathbf{K}_{\Gamma}$ denote the sub-algorithmic complexity $\mathbf{K}_{f_{\Gamma}}$. That is, $\mathbf{K}_{\Gamma}\left(x\right)$ gives the shortest program that can compute or emulate a generative process of the object $x$ in the assembly space $\Gamma$. In the case of infinite assembly spaces, one analogously defines the sub-algorithmic complexity $\mathbf{K}_{\mathcal{S}}$ and the sub-universal a priori probability upon language $\mathbf{L}_{\mathcal{S}}$.
A.2.3 Theoretical results
The main idea to achieve the following theoretical results is to construct a randomly generated program that receives a formal theory (which contains all the computable procedures and statistical criteria in assembly theory) as input. Then, it searches for a molecule (or object in an assembly space) with MA sufficiently high so as to make the pathway probability of spontaneous formation be sufficiently lower than the very own deceiving program’s algorithmic probability, so that the divergence between these two probability distributions become statistically significant according to the arbitrarily chosen statistical method and significance level.
Lemma A.1.
Let $\mathcal{S}$ be infinite computably enumerable. Let $\mathbf{F}$ be an arbitrary formal theory that contains assembly theory, including all the decidable procedures of the chosen method for calculating the assembly index (or approximating MA) of an object for a nested subspace of $\mathcal{S}$, and the program that decides whether or not the criteria for building the assembly spaces are met. Let $k\in\mathbb{N}$ be an arbitrarily large natural number. Then, there are a program $\mathrm{p}_{y}$, $\Gamma\subset\mathcal{S}$ and $y\in V\left(\Gamma\right)$ such that
$\mathbf{K}\left(y\right)+k\leq\left|\mathrm{p}_{y}\right|+k+\mathbf{O}(1)\leq c_{\Gamma}\left(y\right)\text{ ,}$ (3)
where the function $\;c_{\Gamma}\colon\Gamma\subset\mathcal{S}\to\mathbb{N}\;$ gives the MA of the object $y$ in the assembly space $\Gamma$ (or $\mathcal{S}$) and $\mathbf{U}\left(\mathrm{p}_{y}\right)=y$.
Proof.
Let $p$ be a bit string that represents an algorithm running on a prefix universal Turing machine $\mathbf{U}$ that receives $\mathbf{F}$ and $k$ as inputs. Then, it calculates $\left|p\right|+\left|\mathbf{F}\right|+\mathbf{O}\left(\log_{2}\left(k\right)\right)+k$ and enumerates $\mathcal{S}$ while calculating $c_{\Gamma}\left(x\right)$ of the object (or vertex) $x\in V\left(\Gamma\right)\subset V\left(\mathcal{S}\right)$ at each step of this enumeration. Finally, the algorithm returns the first object $y\in V\left(\mathcal{S}\right)$ for which
$\left|p\right|+\left|\mathbf{F}\right|+\mathbf{O}\left(\log_{2}\left(k\right)\right)+k+\mathbf{O}(1)\leq c_{\Gamma}\left(y\right)$ (4)
holds. In order to demonstrate that $p$ always halts, just note that $\mathcal{S}$ is infinite computably enumerable. Also, for any value of $c_{\Gamma^{\prime}}\left(z\right)$ for some $z\in V\left(\Gamma^{\prime}\right)\subset V\left(\mathcal{S}\right)$, there is only a finite number of minimum rooted assembly subspaces (starting on any object in $B_{\mathcal{S}}$ and ending on $z$) whose augmented cardinality is $c_{\Gamma^{\prime}}\left(z\right)$, where $B_{\mathcal{S}}$ is the basis (i.e., the finite set of basic building blocks [cronin]) of $\mathcal{S}$. This implies that there is an infinite number of distinct values of $c_{\Gamma^{\prime}}\left(z\right)$. Now, let $\mathrm{p}_{y}=\left<k,\mathbf{F},p\right>$. Finally, from Equation 4 and basic properties in AIT, we have it that
$\mathbf{K}\left(y\right)+k\leq\left|\mathrm{p}_{y}\right|+k+\mathbf{O}(1)\leq\left|p\right|+\left|\mathbf{F}\right|+\mathbf{O}\left(\log_{2}\left(k\right)\right)+k+\mathbf{O}(1)\leq c_{\Gamma}\left(y\right)$ (5)
holds for some sufficiently large $k$. ∎
Lemma A.2.
Let the conditions for Lemma A.1 be satisfied. Let
$\mathbf{P}\colon\left\{\Gamma^{\prime}\middle|\exists x\left(c_{\Gamma}\left(y\right)=x\right),\,\Gamma^{\prime}\subseteq\Gamma\subset\mathcal{S}\text{ is rooted},\text{ and }y\in V\left(\Gamma^{\prime}\right)\right\}\to\left[0,1\right]$
be an arbitrary probability measure on the set of pathways in $\mathcal{S}$ and $\mathrm{p}_{\mathbf{P}}$ a program that computes a computable function that outputs an upper bound for $\mathbf{P}$ such that for every $\epsilon^{\prime}>0$ and $\Gamma\subset\mathcal{S}$, there are $\Gamma^{\prime}\subset\mathcal{S}$ and $x\in V\left(\Gamma^{\prime}\right)$ with $\Gamma\subseteq\Gamma^{\prime}$ and $\mathbf{P}\left({\Gamma^{*}}_{x}\right)\leq\mathbf{U}\left(\left<{\Gamma^{*}}_{x},\mathrm{p}_{\mathbf{P}}\right>\right)<\epsilon^{\prime}$. Let $01\geq\epsilon>$ 0 be an arbitrary encodable real number. Let $k\in\mathbb{N}$ be an arbitrarily large natural number. Then, there are a program $\mathrm{p}_{\epsilon}$, $\Gamma\subset\mathcal{S}$, and $y\in V\left(\Gamma\right)$ such that Lemma A.1 is satisfied with $y$ and
$\mathbf{K}\left(y\right)+k\leq\left|\mathrm{p}_{\epsilon}\right|+k+\mathbf{O}(1)\leq c_{\Gamma}\left(y\right)\text{ ,}$ (6)
and
$\mathbf{P}\left({\Gamma^{*}}_{y}\right)\leq\mathbf{U}\left(\left<{\Gamma^{*}}_{y},\mathrm{p}_{\mathbf{P}}\right>\right)<\epsilon$ (7)
hold, where $\mathbf{U}\left(\mathrm{p}_{\epsilon}\right)={\Gamma^{*}}_{y}$.
Proof.
Let $p^{\prime}$ be a bit string that represents an algorithm running on a prefix universal Turing machine $\mathbf{U}$ that receives $\mathrm{p}_{\mathbf{P}}$, $\epsilon$ and $k$ as inputs. Then, it enumerates the assembly pathways ${\Gamma^{*}}_{x}$ in $\mathcal{S}$ such that $\left|p\right|+\left|\mathbf{F}\right|+\mathbf{O}\left(\log_{2}\left(k\right)\right)+\left|p^{\prime}\right|+\left|\mathrm{p}_{\mathbf{P}}\right|+\mathbf{O}\left(\log_{2}\left(\epsilon\right)\right)+k+\mathbf{O}(1)\leq c_{\Gamma}\left(x\right)$, $\mathbf{U}\left(\left<{\Gamma^{*}}_{x},\mathrm{p}_{\mathbf{P}}\right>\right)<\epsilon$, and Lemma A.1 holds for $x$ given $k$. Finally, it returns this first ${\Gamma^{*}}_{x}$ in this enumeration. Now, let $\mathrm{p}_{\epsilon}=\left<\mathrm{p}_{\mathbf{P}},\epsilon,k,p^{\prime}\right>$. Therefore, in addition to the arguments in the proof of Lemma A.1, the desired theorem follows from the fact that program $p^{\prime}$ always halts because of our initial assumptions on program $\mathrm{p}_{\mathbf{P}}$ and the probability distribution given by $\mathbf{P}$. ∎
Lemma A.3.
Let the conditions for Lemmas A.1 and A.2 be satisfied. Let $\mathbf{F}^{\prime}\supseteq\mathbf{F}$ be a formal theory that also includes the chosen statistical method, the criteria for the arbitrarily chosen statistical significance level, the chosen computable method for approximating $\mathbf{P}$ from above with program $\mathrm{p}_{\mathbf{P}}$, and the acceptable maximum error $\mathscr{E}\in\mathbb{N}$ for an overestimation of the complexity of an object in $\mathcal{S}$. Then, there are a program $\mathrm{p}_{d}$, $\Gamma\subset\mathcal{S}$, and $y\in V\left(\Gamma\right)$ such that Lemma A.2 is satisfied with $y$ and $\mathbf{F}^{\prime}$ decides that the divergence of the (sub-)universal distribution from $\mathbf{P}$ is statistically significant, where $\mathbf{U}\left(\mathrm{p}_{d}\right)={\Gamma^{*}}_{y}$ and $\left|\mathrm{p}_{d}\right|+\mathscr{E}<c_{\Gamma}\left(y\right)$ hold.
Proof.
Let $\mathrm{p}_{d}$ be a bit string that represents an algorithm running on a prefix universal Turing machine $\mathbf{U}$ that includes the computation of $\mathrm{p}_{k}$ and $\mathrm{p}_{\epsilon}$ (which are programs defined in the proofs of Lemmas A.1 and A.2) as subroutines. Then, it searches for the first ${\Gamma^{*}}_{x}$, sufficiently small value of $\epsilon$, and sufficiently large value of $k\geq\mathscr{E}+\mathbf{O}(1)$ such that Lemmas A.1 and A.2 are satisfied with $x$, $\left|\mathrm{p}_{d}\right|+\mathscr{E}<\left|p\right|+\left|\mathbf{F}^{\prime}\right|+\mathbf{O}\left(\log_{2}\left(k\right)\right)+\left|p^{\prime}\right|+\left|\mathrm{p}_{\mathbf{P}}\right|+\mathbf{O}\left(\log_{2}\left(\epsilon\right)\right)+k+\mathbf{O}(1)<c_{\Gamma}\left(x\right)$ holds, and the divergence of $2^{-\left|\mathrm{p}_{d}\right|}$ from $\mathbf{U}\left(\left<{\Gamma^{*}}_{x},\mathrm{p}_{\mathbf{P}}\right>\right)$ is statistically significant according to the formal theory $\mathbf{F}^{\prime}$. Finally, the algorithm returns this first assembly pathway ${\Gamma^{*}}_{x}$ as output. Note that, since the value of $2^{-\left|\mathrm{p}_{d}\right|}$ is fixed, one can always employ program $\mathrm{p}_{\mathbf{P}}$ and the statistical criteria in theory $\mathbf{F}^{\prime}$ to find an arbitrarily lower probability than $2^{-\left|\mathrm{p}_{d}\right|}$ so that the resulting probability distribution (i.e., the probability distribution given by $\mathbf{P}$) diverges from the sub-universal (a priori probability) distribution. This holds because: of the algorithmic coding theorem, which implies that $2^{-\left|\mathrm{p}_{d}\right|}$ is a lower bound for the sub-universal a priori probability upon language $\mathbf{L}^{\prime}$, where $\mathrm{p}_{d}\in\mathbf{L}^{\prime}\subseteq\mathbf{L}$; and of the fact that $\mathrm{p}_{\mathbf{P}}$ is a program that computes a (computable) function that outputs an upper bound for $\mathbf{P}$888 Also note that $\mathbf{F}^{\prime}$ does not actually need to be able to compute the value of the sub-universal a priori probability of $x$ because one already knows $2^{-\left|\mathrm{p}_{d}\right|}$ is a lower bound for it and $\mathbf{U}\left(\left<{\Gamma^{*}}_{x},\mathrm{p}_{\mathbf{P}}\right>\right)$ is an upper bound for the optimal pathway probability of $x$. Additionally, this divergence eventually becomes statistically significant (as the divergence increases) because the probability of occurrence of a sequence of events following an empirical probability distribution, which diverges from the original distribution that the events are generated, eventually decreases as the divergence sufficiently increases. Also note that $\left|\mathrm{p}_{d}\right|\leq\left|\mathrm{p}_{k}\right|+\left|\mathrm{p}_{\epsilon}\right|+\mathbf{O}(1)$. Therefore, since $k$ and $\epsilon$ were arbitrary in Lemma A.2 and all the statistical methods in $\mathbf{F}^{\prime}$ are decidable by assumption, we have that $\mathrm{p}_{d}$ always halts. ∎
Theorem A.4.
Let the conditions for Lemma A.3 be satisfied. Let $\mathcal{S}$ be an infinite assembly space whose set of randomly generated (computable) generative processes include one that can effect at least the computation of program $\mathrm{p}_{d}$, where Lemma A.3 holds for $\mathrm{p}_{d}$ and $y\in V\left(\mathcal{S}\right)$. Then:
  • the complexity error is larger than $\mathscr{E}$ (except for an independent constant);
  • and this error implies a statistically significant (according to $\mathbf{F}^{\prime}$) distinct frequency of occurrence of $y$ than it was expected from the chosen assembly theory.
Proof.
From Lemma A.3, we have it that $\mathbf{U}\left(\mathrm{p}_{d}\right)={\Gamma^{*}}_{y}$. Thus, from our assumptions and the definition of $\mathbf{K}_{\mathcal{S}}$, we have it that $\mathbf{K}_{\mathcal{S}}\left(y\right)\leq\left|\mathrm{p}_{d}\right|+\mathbf{O}(1)$, which proves that the complexity error is larger than $\mathscr{E}$ from Lemma A.3. We also have that the probability of an assembly pathway being constructed by a randomly generated computable process is given by the sub-universal a priori probability of ${\Gamma^{*}}_{y}$ upon language $\mathbf{L}_{\mathcal{S}}$, i.e.,
$\sum\limits_{\small\begin{array}[]{c}\mathbf{U}\left(p\right)={\Gamma^{*}}_{y};\\ p\in\mathbf{L}_{\mathcal{S}}\end{array}}C\,\frac{1}{2^{\left|p\right|}}\text{ .}$ (8)
Therefore, by replacing $\mathbf{L}^{\prime}$ with $\mathbf{L}_{\mathcal{S}}$ in the proof of Lemma A.3, we achieve a statistically significant (according to $\mathbf{F}^{\prime}$) distinct frequency of occurrence of $y$ than it was expected from the chosen assembly theory. ∎
Corollary A.5.
Let the conditions for Lemma A.3 be satisfied. Let $\mathcal{S}$ be an infinite assembly space whose randomly generated (computable) generative processes are capable of universal computation. Then:
  • the complexity error is larger than $\mathscr{E}$ (except for an independent constant);
  • and this error implies a statistically significant (according to $\mathbf{F}^{\prime}$) distinct frequency of occurrence of $y$ than it was expected from the chosen assembly theory.
Proof.
The proof follows directly from the fact that $\mathbf{K}\left(y\right)\leq\mathbf{K}_{f}\left(y\right)+\mathbf{O}(1)$ and from replacing $\mathbf{L}^{\prime}$ with $\mathbf{L}$ in the proof of Lemma A.3. ∎
A.3 PyBDM Code for CTM and BDM
The Coding Theorem (CTM) and Block Decomposition Methods (BDM) are resource-bounded computable methods [soler, bdm, Zenil2020] that attempt to approximate semi-computable measures that are a generalisation of statistical measures more powerful than the methods proposed in “Assembly Theory” as they combine global calculations of classical entropy with local estimations of algorithmic information content.
Algorithm 1 Python implementation of 2D-Block Decomposition Method (PyBDM)
import numpy as np
from pybdm import BDM
import pandas as pd
X = pd.read_csv(r’file directory’,dtype=int)
bdm = BDM(ndim=2)
Z=X.to_numpy()
bdm.bdm(Z)